### v2

@@ Abstract
Transformers come with a high computational cost, yet their effectiveness in addressing problems in language and vision has sparked extensive research aimed at enhancing their efficiency. However, diverse experimental conditions, spanning multiple input domains, prevent a fair comparison based solely on reported results, posing challenges for model selection. To address this gap in comparability, we design a comprehensive benchmark of more than 30 models for image classification, evaluating key efficiency aspects, including accuracy, speed, and memory usage. This benchmark provides a standardized baseline across the landscape of efficiency-oriented transformers and our framework of analysis, based on Pareto optimality, reveals surprising insights. Despite claims of other models being more efficient, ViT remains Pareto optimal across multiple metrics. We observe that hybrid attention-CNN models exhibit remarkable inference memory- and parameter-efficiency. Moreover, our benchmark shows that using a larger model in general is more efficient than using higher resolution images. Thanks to our holistic evaluation, we provide a centralized resource for practitioners and researchers, facilitating informed decisions when selecting transformers or measuring progress of the development of efficient transformers.  [^1] https://github.com/tobna/WhatTransformerToFavor
Keywords:
Vision Transformer Efficiency Efficient Transformer
@@ 1 Introduction
Figure 1: Pareto front (dotted line) of throughput and accuracy. Marker shapes and hues signify the different overarching strategies for making ViT more efficient. Pareto optimal models have a black dot, while the others have a white dot. The size of markers signifies if we fine-tuned using the standard $224$px or the larger $384$px resolution. We mark the baselines ViT-Ti@224 (A), ViT-Ti@384 (B), and ViT-S@224 (C) in the plot.
@@
The introduction of the Transformer [55] prompted its widespread adoption in both language and vision communities. In particular for image classification, the Vision Transformer (ViT) [14] has positioned itself as one of the best known applications of the original architecture, outperforming traditional CNN architectures on benchmarks like ImageNet [68, 67]. However, a major challenge in working with Transformer models is dealing with the computational complexity of the self-attention mechanism. This mechanism enables the Transformer to capture global dependencies between pairs of sequence positions, but it has a computational complexity of $\mathcal{O}(N^{2})$ in the input length $N$, making it impractical for long sequences and high-resolution images.
@@
Efforts have been made to reduce the computational load of Transformer models, particularly in resource-constrained settings, such as embedded systems [56, 44]. Researchers have explored numerous strategies, like implementing sparse, local, or kernelized attention mechanisms, as well as token removal criteria to decrease the sequence length. However, selecting the most efficient model that meets certain performance standards remains a challenging task. This task becomes particularly difficult, knowing that “efficiency” can refer to different concepts, like training resources, inference speed, memory requirements, or number of parameters. Moreover, difficulties identifying the most efficient, yet highly performant vision transformers are accentuated by different training and evaluation conditions reported in the literature. It is even more unclear how tradeoffs in efficiency for models proposed for NLP carry over to vision tasks.
@@
To address these problems of comparing recent advances in efficient ViTs, we design an impartial test bed, training models from scratch on the same data, using the same setup. This enables the measurement of multiple real-world performance metrics under consistent conditions. We identify common strategies employed to enhance model efficiency and utilize them to correlate changes in efficiency metrics with the respective strategies. Through this test bed, we provide a centralized resource of comprehensive baselines in the domain of image classification, conducting a thorough review of the current state of research on efficient Transformer models. While many papers at most report speed [35, 11, 32], we provide empirical values across multiple dimensions of efficiency using the same setup. In particular, we also provide baselines for efficient Transformers, that have been proposed for NLP, but can be trivially adapted to CV.
@@
We analyze model speed and memory requirements both at training and inference time, as well as parameter efficiency. This reveals that not all efficiency claims are realized. Our analysis is based on the Pareto front, the set of models that provide an optimal tradeoff between model performance and one aspect of efficiency. We plot inference throughput against accuracy in Figure 1 to visualize the Pareto front. Here, a model is Pareto optimal (black dot) if and only if there is no model that is both more accurate and faster at the same time. Through this analysis, we show that ViT remains Pareto optimal across three of the four empirical metrics we track. But we also find other efficiency strategies that are Pareto optimal for different metrics. This benchmark, based on more than 150 experiments, provides researchers and practitioners with a valuable resource, aiding them in choosing the most efficient and effective model architecture for their specific use case.
@@
  • A comprehensive one-stop benchmark of transformer-based models. Experiments are run under the same conditions, making the results comparable.
  • We find that in most cases, scaling up the model is more efficient, being faster and more accurate, than scaling up the image resolution.
  • We find that ViT still is Pareto optimal for three out of four metrics; in particular, for training and inference speed. We provide more detailed suggestions on which models to use, depending on the main constraint, in Figure 6.
@@ 2 Related Work
This section provides an overview of relevant work on transformer surveys and approaches to measure efficiency, while Section 3.3 delves into specific efficient transformer-like architectures.
@@
Efficiency has become a critical aspect of transformers, leading to the exploration of various strategies and evaluations across domains. Surveys on efficient transformers provide valuable taxonomies and insights. [48] focus on efficiency gains primarily in NLP, while [17] and [70] collect general approaches for enhancing model efficiency, including those applicable to transformers. For ViTs, [35] presents an extensive list of efficient versions, classified based on different aspects such as computational complexity, robustness, and transparency. They, along with [30], compare the efficiency of ViTs in terms of parameters, ImageNet accuracy, and other performance metrics using data from the original papers. Surveys on ViT-like models, conducted by [18], [64], and [22], focus on categorizing models according to different vision tasks, while [71] specializes in dense prediction. Specialized surveys investigate the application of transformers in specific domains, such as action recognition [53], image restoration [2], medical imaging [43, 27, 20, 34], or remote sensing [1], or using other modalities like speech recognition [25], language processing [7], time series analysis [58], or multimodal tasks [61].
In contrast, we survey efficient Transformers as a first step to be able to evaluate their efficiency empirically under the same conditions using more than 150 experiments.
@@
Efficiency evaluation in deep learning models is another area of investigation. [4] provides a theoretical overview of efficiency aspects and metrics, along with measurement methodologies. [12] offers a detailed discussion on efficiency metrics, highlighting the potential pitfalls of relying solely on theoretical metrics. [6] conducts a highly regarded survey on the efficiency of CNNs, and [29] compares the efficiency of their novel architectures with older models using the Pareto front of throughput and accuracy. Finally, [47] constructs a comprehensive benchmark on synthetic data to quantify various aspects of transformer model performance, facilitating efficiency evaluations.
We build on this work and contribute by evaluating more than 30 efficient Transformers for CV using real data. We analyze our benchmark’s results in-depth.
@@ 3 Efficient Transformers for Vision
First, we present our methodology for efficiency assessment, before briefly describing ViT, as well as a diverse collection of efficient transformers that have been proposed and what these change compared to ViT. These models, from the domains of CV and NLP, all claim to improve the efficiency of the baseline transformer. We systematically select efficient transformers based on diversity, popularity, and novelty of their approaches.
@@ 3.1 Quantifying Efficiency
The term efficiency can take on different meanings; it is therefore crucial to consider multiple dimensions when evaluating a model’s efficiency. In this paper, we analyze model throughput in images per second and memory requirements in GB of VRAM at training and inference time and contrast them with accuracy.
Due to the complex multidimensional tradeoffs, efficiency cannot be captured by a single number. Hence, we focus on the Pareto front to identify models that achieve the best trade-offs between two metrics. The Pareto front represents the set of models that take the best compromises between two metrics. A Pareto optimal model outperforms every other model in at least one of the two metrics. See the Pareto front in Figure 1, where for each model that is not on the Pareto front (white dot) there is a model on the Pareto front (black dot) that is both faster and more accurate.
@@ 3.3 Efficiency-Improving Changes
After having gone over the backbone that constitutes a ViT, we can discuss the most important modifications, which have been proposed to make it more efficient. In order to analyze the overarching strategies that are taken, we classify the efficient models systematically using a two-step approach. First, we identify the specific components – the token mixing mechanism, the token sequence, or the MLP block – of the baseline ViT architecture that are modified (i.e., “where”), as visualized in Figure 2. On the second level, the taxonomy is structured based on the strategies to enhance the efficiency (i.e., “how”). While this taxonomy is not meant to be a comprehensive overview of ViT-like models, it is proposed as a tool for identifying the most popular strategies to make vision transformers more efficient. Note that our categorization scheme aligns closely with taxonomies that have been recently proposed [17, 48, 41].
3.3.1 (i) Token Mixing
The first and most popular approach is to change the token mixing mechanism, which directly tackles the $\mathcal{O}(N^{2})$ computational complexity of self-attention. There are multiple strategies through which this can be accomplished: Some methods approximate the attention mechanism with reduced computation, which can be achieved by matrix decomposition, changing the order of operations, or fixing attention values. Other approaches combine attention with CNNs to perform sub-sampling in the attention mechanism or reduce the number of uses of the attention mechanism. Finally, some methods discard the attention mechanism and instead introduce entirely new token mixing strategies.
@@
Here, $Q^{\top}K\in\mathbb{R}^{d\times d}$ is used to replace the low-rank matrix $QK^{\top}$, to define the globally informed filter ${\operatorname{softmax}\left(\frac{1}{\sqrt{d}}Q^{\top}K\right)\in\mathbb{R}^{d\times d}}$ which is applied to each token individually. Since both $Q$ and $K$ are likely to be of rank $d$, the former most likely has full rank, which enables more efficient information encoding.
Sparse attention is an alternative approach to these dynamic approximations of global interactions. Normal attention tends to focus on only a few input tokens, giving low attention scores to most other tokens [23]. Sparse attention takes advantage of this by setting most of the attention weights to zero and only calculating the most important ones. The parts of the attention matrix, which are determined dynamically, can be based on a fixed pattern, like in the widely used Swin Transformer [32], and SwinV2 [31], where attention is performed only inside local sets of image tokens, or in HaloNet [54], where each token can only attend to its local neighborhood. Alternatively, Routing Transformer [39] determines sets of tokens that exchange information by grouping them based on their content. In contrast, Sinkhorn Transformer [46] fixes local groups and lets tokens only attend to a different group using a dynamic permutation matrix. Another way of keeping the attention matrix sparse, is by sub-sampling the keys and values. For example, Wave-ViT [65] uses a discrete wavelet transform for sub-sampling.
Fixed attention is an extreme example of setting attention values beforehand. In this case, attention values only depend on a tokens position. This is explored in the Synthesizer [45].
@@
which can be calculated with linear complexity $\mathcal{O}(N)$ in the number of tokens. Various kernels have been proposed, including the random Gaussian kernel, introduced with the Performer [11] and expanded on in the FourierLearner-Transformer [10] by adding a relative positional encoding in phase space. Scatterbrain [9] combines the Performers’ attention mechanism with sparse attention. Poly-SA [3] uses the identity function $\varphi\equiv\mathrm{id}$, Linear Transformer [21] uses an ELU, and another variant of the Performer uses a ReLU kernel.
@@
Hybrid attention approaches combine convolutions with the attention mechanism. EfficientFormer(-V2) [28] uses convolutions initially to focus on local interactions, and then employs the attention mechanism to capture global interactions. Another approach is to use convolutions inside the attention mechanism to create locally informed queries, keys, and values, as in CvT [59], where it is used to subsample keys and values, and in CoaT [62].
Non-attention shuffling refers to a group of techniques to capture global interactions without using an attention mechanism. MLP-Mixer [49] applies a fully connected layer along the token sequence for global interactions, while FocalNet [63] extracts a hierarchy of contexts using convolutions. Moreover, FNet [26, 42] directly exploits the Fast Fourier Transform (FFT) for $\mathcal{O}(N\log N)$ complexity token mixing, while GFNet [37] utilizes it to implement a global convolution.
3.3.2 (ii) Token Sequence
@@
The second category, token sequence, is more prevalent in efficient transformers used in computer vision compared to natural language processing. The idea here is to remove redundant information typically contained in images, and in doing so, reducing computational costs without significantly affecting the model’s performance. Methods aim to reduce the token sequence by removing unnecessary image patches, such as those associated with the background, by merging similar patches to minimize redundant information, or by summarizing the information into a smaller number of abstract tokens to represent higher-level information. These models leverage the $\mathcal{O}(N^{2})$ complexity of the self-attention mechanism to attain a large reduction in computational cost, as removing $30\%$ of the tokens reduces the operations needed by around $50\%$. While some of these methods can directly be applied to a trained model, performance improves when utilizing them during training [5], like we do.
@@
Token Removal requires identifying which tokens to remove while without losing critical information. To that end, Dynamic ViT [36] uses the Gumbel-softmax for token retention probabilities, and A-ViT [66] learns halting probabilities that weight the token outputs of different depths. In contrast, EViT [29] avoids introducing extra parameters by utilizing the previous layer’s attention matrix to remove tokens.
@@
Token merging is another strategy to remove redundant information While a version of EViT [29] merges the unimportant tokens, ToMe [5] merges tokens based on their similarity, using a fast bipartite matching algorithm.
@@
3.3.3 (iii) MLP Block
The final way in which proposed methods change the architecture of transformers is by moving compute to the MLP block, which has linear complexity with respect to the sequence length. Despite room for efficiency gains, we have identified just one method taking this approach, which we include to make this benchmark more diverse. Switch Transformer [16] focuses on improving performance while maintaining a similar computational cost.
Specialized parameters can be used to tackle this problem by introducing multiple parameter sets for each MLP block and passing different tokens through blocks with different parameters. Switch Transformer can have multiple times the number of parameters of the ViT without introducing lots of extra computations.
To recap, the categorization we present offers a structured framework for understanding the diverse approaches to improve the efficiency of ViT-like models. This enables us to compare broad trends across strategies.

### v3

@@ Abstract
Self-attention in Transformers comes with a high computational cost because of their quadratic computational complexity, but their effectiveness in addressing problems in language and vision has sparked extensive research aimed at enhancing their efficiency. However, diverse experimental conditions, spanning multiple input domains, prevent a fair comparison based solely on reported results, posing challenges for model selection. To address this gap in comparability, we perform a large-scale benchmark of more than 45 models for image classification, evaluating key efficiency aspects, including accuracy, speed, and memory usage. Our benchmark provides a standardized baseline for efficiency-oriented transformers. We analyze the results based on the Pareto front – the boundary of optimal models. Surprisingly, despite claims of other models being more efficient, ViT remains Pareto optimal across multiple metrics. We observe that hybrid attention-CNN models exhibit remarkable inference memory- and parameter-efficiency. Moreover, our benchmark shows that using a larger model in general is more efficient than using higher resolution images. Thanks to our holistic evaluation, we provide a centralized resource for practitioners and researchers, facilitating informed decisions when selecting or developing efficient transformers.  [^1] https://github.com/tobna/WhatTransformerToFavor
@@ 1 Introduction
Figure 1: Pareto front (dotted line) of throughput and accuracy. Marker shapes and hues signify the different overarching strategies for making ViT more efficient. Pareto optimal models have a black dot, while the others have a white dot. The size of markers signifies if we fine-tuned using the standard $224$px or the larger $384$px resolution. We mark the baselines ViT-Ti@224 (A), ViT-Ti@384 (B), and ViT-S@224 (C) and also plot the Pareto front for high resolution images (dashed line).
@@
The introduction of the Transformer [63] prompted its widespread adoption in both language and vision communities. In particular for image classification, the Vision Transformer (ViT) [16] has positioned itself as one of the best known applications of the original architecture, outperforming traditional CNN architectures on benchmarks like ImageNet [76, 75]. However, a major challenge in working with Transformer models is dealing with the computational complexity of the self-attention mechanism. This mechanism enables the Transformer to capture global dependencies between pairs of tokens, but it has a computational complexity of $\mathcal{O}(N^{2})$ in the input length $N$, making it impractical for long sequences and high-resolution images.
@@
Efforts have been made to reduce the computational load of Transformer models, particularly for resource-constrained settings, such as embedded systems [64, 51]. Researchers have explored numerous strategies, like implementing sparse, local, or kernelized attention mechanisms, as well as token removal criteria to decrease the sequence length. However, selecting the most efficient model that meets certain performance standards remains a challenging task. This task becomes particularly difficult, knowing that “efficiency” can refer to different concepts, like training resources, inference speed, or memory requirements. Moreover, difficulties identifying the most efficient, yet highly performant vision transformers are accentuated by different training and evaluation conditions reported in the literature. It is even more unclear how tradeoffs in efficiency for models proposed for NLP carry over to vision tasks.
@@
To address these problems of fairly comparing efficient ViTs, we design an impartial test bed, training models from scratch on the same data, using the same setup. While many papers at most report inference speed [41, 37, 38, 61], our benchmark enables us to fairly provide empirical values across multiple dimensions of efficiency. We identify common strategies employed to enhance model efficiency and utilize them to correlate changes in efficiency metrics with the respective strategies. Through our benchmark, we provide a centralized resource of comprehensive baselines in the domain of image classification, conducting a thorough review of the current state of research on efficient Transformers. In particular, we provide baselines for Transformers proposed for NLP, which can be trivially adapted to CV.
@@
We analyze model speed and memory requirements both at training and inference time, as well as parameter efficiency. This reveals that not all efficiency claims are realized. Our analysis is based on the Pareto front, the set of models that provide an optimal tradeoff between model performance and one aspect of efficiency. It lets us analyze the complex multidimensional tradeoffs involved in judging efficiency. We plot inference throughput against accuracy in Figure 1 to visualize the Pareto front. Here, a model is Pareto optimal (black dot) if and only if there is no model that is both more accurate and faster at the same time. Through this analysis, we show that ViT remains Pareto optimal across three of the four empirical metrics we track, and also find other efficiency strategies that are Pareto optimal for different metrics. This benchmark, based on more than 200 experiments, provides researchers and practitioners with a valuable resource, aiding them in choosing the most efficient and effective model architecture for their specific use case.
@@
  • A fair large-scale benchmark for efficiency in transformer models run on 45+ models, closing the comparability gap due to different experimental conditions.
  • We find that a well-trained ViT still is Pareto optimal for three out of four metrics, challenging claims of other models being more efficient [68, 43, 10].
  • Our analysis uncovers novel findings like that in most cases, scaling up the model size is more efficient than scaling up the image resolution.
@@ 2 Related Work
This section provides an overview of relevant work on transformer surveys and approaches to measure efficiency, while Section 3.3 delves into specific efficient architectures.
@@
Efficiency has become a critical aspect of transformers, leading to the creation of multiple surveys on efficient transformers providing valuable taxonomies and insights. [55] focuses on efficiency gains primarily in NLP, while [19] and [81] collect general approaches for enhancing model efficiency, including those applicable to transformers. [41, 35] present an extensive list of efficient ViTs, classified based on aspects such as computational complexity, robustness, and transparency, comparing the efficiency in terms of parameters, ImageNet accuracy, and inference speed, using information from the original papers. Surveys on ViT-like models [22, 72, 26] focus on categorizing models according to different vision tasks, while [82] specializes in dense prediction. Specialized surveys investigate the application of transformers in specific domains, like action recognition [60], image restoration [2], medical imaging [50, 31, 23, 40], remote sensing [1], speech recognition [29], language processing [9], time series [66], or multimodal tasks [69].
We build on this by analyzing efficiency gains in terms of broader strategies for efficiency. In contrast to previous work, we close the gap of different training strategies and enable a fair comparison of efficient architectures under the same conditions based on more than new 200 experiments.
@@
Efficiency evaluation in deep learning models is another area of investigation. [4, 14] provide an overview of efficiency aspects and metrics, along with measurement methodologies, highlighting the potential pitfalls of relying solely on theoretical metrics. [8] conducts a highly regarded survey on the efficiency of CNNs, and [34] compares the efficiency of their novel architectures with older models using the Pareto front of throughput and accuracy. Finally, [54] introduces a benchmark on synthetic data to quantify transformer model performance on long sequences in NLP, evaluating 10 models.
We extend previous work by evaluating more than 45 efficient Transformers for CV, measuring empirical metrics on the same hardware and using real data. The in-depth analysis of our benchmark’s results reveals new insights.
@@ 3 Efficient Transformers for Vision
First, we present our methodology for efficiency assessment. Then we briefly describe ViT, as well as a diverse collection of efficient transformers that have been proposed and what these change compared to ViT. These models from the domains of CV and NLP all claim to improve the efficiency of the baseline transformer. We systematically select efficient transformers based on popularity, diversity, and novelty of their approaches. See the supplementary material for more detailes on model selection.
@@ 3.1 Quantifying Efficiency
The term efficiency can take on different meanings; it is therefore crucial to consider multiple dimensions when evaluating a model’s efficiency. In this paper, we analyze model speed and memory requirements at training and inference time and contrast them with accuracy.
Due to the complex multidimensional tradeoffs, efficiency cannot be captured by a single number. Hence, we focus on the Pareto front – the boundary of the landscape of efficient models – to identify models that achieve the best trade-offs between two metrics. The Pareto front represents the set of models that take the best compromises between two metrics. A Pareto optimal model outperforms every other model in at least one of the two metrics. See the Pareto front in Figure 1, where for each model that is not on the Pareto front (white dot) there is a model on the Pareto front (black dot) that is both faster and more accurate.
@@ 3.3 Efficiency-Improving Changes
After having gone over the backbone that constitutes a ViT, we discuss the most important modifications, which have been proposed to make it more efficient. In order to analyze the overarching strategies that are taken, we classify the efficient models systematically using a two-step approach. First, we identify the specific components — the token mixing mechanism, the token sequence, or the MLP block — of the baseline ViT architecture that are modified (i.e., “where”), as visualized in Figure 2. On the second level, we classify the different strategies to enhance the efficiency (i.e., “how”). While this taxonomy is not meant to be a comprehensive overview of ViT-like models, it is proposed as a tool for identifying the best and most popular strategies to make vision transformers more efficient. Note that our categorization scheme aligns closely with taxonomies that have been recently proposed [19, 55, 47].
(i) Token Mixing
The first and most popular approach is to change the token mixing mechanism, which directly tackles the $\mathcal{O}(N^{2})$ computational complexity of self-attention. There are multiple strategies through which this can be accomplished. Some methods approximate the attention mechanism with reduced computation, which can be achieved by matrix decomposition, changing the order of operations, or fixing attention values. Other approaches combine attention with CNNs to perform sub-sampling in the attention mechanism or reduce the number of uses of the attention mechanism. Finally, some methods discard the attention mechanism and instead introduce entirely new token mixing strategies.
@@
Here, $Q^{\top}K\in\mathbb{R}^{d\times d}$ is used to replace the low-rank matrix $QK^{\top}$ to define the globally informed filter ${\operatorname{softmax}\left(\frac{1}{\sqrt{d}}Q^{\top}K\right)\in\mathbb{R}^{d\times d}}$ which is applied to each token individually. Since both $Q$ and $K$ are likely to be of rank $d$, the former most likely has full rank, which enables more efficient information encoding.
Sparse attention is an alternative approach to these dynamic approximations of global interactions. Normal attention tends to focus on only a few input tokens, giving low scores to most others [27]. Sparse attention takes advantage of this by fixing most of the attention weights at zero and only calculating the most important ones. The parts of the attention matrix, which are determined dynamically, can be based on a fixed pattern, like in the widely used Swin Transformer [37], and SwinV2 [36], where attention is performed only inside local sets of image tokens, or in HaloNet [62], where each token can only attend to its neighbors. Alternatively, Routing Transformer [45] determines sets of tokens that exchange information by grouping them based on their content. In contrast, Sinkhorn Transformer [53] fixes local groups and lets tokens only attend to a different group using a dynamic permutation matrix. Informer [80] only calculates attention values for the $c\cdot\ln N$ most important tokens. Another way of keeping the attention matrix sparse, is by sub-sampling the keys and values. For example, Wave-ViT [73] uses a discrete wavelet transform for sub-sampling.
Fixed attention is an extreme example of setting attention values beforehand, which then only depend on the tokens positions. This is explored in the Synthesizer [52].
@@
which can be calculated with linear complexity $\mathcal{O}(N)$ in the number of tokens. Various kernels have been proposed, including the random Gaussian kernel, introduced in Performer [13] and expanded on in the FourierLearner-Transformer [12] by adding a relative positional encoding in phase space. Scatterbrain [11] combines this attention mechanism with sparse attention. Poly-SA [3] uses the identity function $\varphi\equiv\mathrm{id}$, Linear Transformer [25] uses an ELU, and SLAB [21] uses a ReLU kernel. Hydra ViT [6] uses an L2-normalization kernel and employs $H=d$ heads.
@@
Hybrid attention combines convolutions with the attention mechanism. EfficientFormerV2 [33] uses convolutions initially to focus on local interactions, and then employs the attention mechanism to capture global interactions, while EfficientViT [7] utilizes convolutions first and kernel attention, using the ReLU kernel later on. Next-ViT [32] alternates $N_{\text{Conv}}\in\mathbb{N}$ convolution blocks with a single attention block. Another approach is to use convolutions inside the attention mechanism to create locally informed queries, keys, and values, as in CvT [67] and ResT [77], where it is used to subsample keys and values, and in CoaT [70].
Fourier attention uses the Fast Fourier Transform (FFT) in the calculation of the attention mechanism to reduce the complexity to $\mathcal{O}(N\log N)$. FNet [30, 48] directly exploits the FFT for $\mathcal{O}(N\log N)$ complexity token mixing, while GFNet [43] utilizes it for a global convolution and AFNO [20] uses an MLP in Fourier space.
Non-attention shuffling refers to techniques of capturing token interactions without using attention. For example MLP-Mixer [56] uses a fully connected layer for global interactions, while FastViT [61] mixes tokens using depth-wise convolutions. EfficientMod [38] modulates a convolutional context with a value matrix and FocalNet [71] modulates a hierarchy of contexts extracted by convolutions. SwiftFormer [49] calculates a global query to multiply with the keys and also replaces the MLP block by a CNN.
(ii) Token Sequence
@@
The second category, token sequence, is more prevalent in efficient transformers used in CV compared to NLP. The idea is to remove redundant information typically contained in images, and in doing so, reducing computational costs without significantly affecting the model’s performance. Methods aim to reduce the token sequence by removing unnecessary image patches, such as the background, by merging tokens to minimize redundant information, or by summarizing the information into a smaller number of abstract tokens representing higher-level information. These models leverage the $\mathcal{O}(N^{2})$ complexity of the self-attention mechanism to attain a large reduction in computational cost, as removing $30\%$ of the tokens reduces the operations needed by around $50\%$. While some of these methods can directly be applied to trained models, performance improves when implementing them during training [5], which we do.
@@
Token Removal requires identifying which tokens to remove without losing critical information. To that end, Dynamic ViT [42] uses the Gumbel-softmax for token retention probabilities, and A-ViT [74] learns halting probabilities that weight the token outputs of different depths. In contrast, EViT [34] avoids introducing extra parameters by utilizing the previous layer’s attention matrix.
@@
Token merging is another strategy to remove redundant information. While a version of EViT [34] merges the unimportant tokens, ToMe [5] merges tokens based on their similarity, using a fast bipartite matching algorithm.
@@
(iii) MLP Block
The final way in which proposed methods change the architecture of transformers is by moving compute to the MLP block, which has linear complexity with respect to the sequence length. Despite room for efficiency gains, we have identified only two methods taking this approach, which we include to make this benchmark more diverse.
More MLPs can be used to tackle this problem by introducing multiple parameter sets for each MLP block and passing different tokens through blocks with different parameters. Switch Transformer [18] can have multiple times the number of parameters of the ViT without introducing lots of extra computations. In contrast, HiViT [78] exchanges early attention layers for additional MLP blocks.
Out categorization offers a structured framework for understanding approaches of improving ViT efficiency. This enables us to compare broad trends across strategies.